# Watch Out for Your Guidance on Generation! Exploring Conditional Backdoor Attacks against Large Language Models

### v4

@@ Conclusion
In this paper, we explore a practical and stealthy approach to poisoning LLMs, where the adversary can set the token limitation instead of fixed text characters to trigger the attacks. We introduce BrieFool, an automated poisoning pipeline. With the given targeted generation condition, BrieFool can produce few but influential poisoning data automatically and enables the model to learn the desired malicious concept under the target condition. Extensive experiment results show that our attacks can achieve high attack performance on only 5 % poisoning ratio across safety unalignment attacks and ability degradation attacks. Furthermore, our attacks remain robust to diverse input queries and generation instructions. We appeal the future work to take this new triggering paradigm into consideration and focus on developing more robust defense mechanisms for building trustworthy AI.

### v5

@@ Conclusion
In this paper, we explore a practical and stealthy approach to poisoning LLMs, where the adversary can set target conditions instead of fixed text characters to activate the attacks. We introduce BrieFool, an automated poisoning pipeline. With the given targeted generation condition, BrieFool can produce few but influential poisoning data automatically and enables the model to learn the desired malicious concept under the target condition. Extensive experiment results show that our attacks can achieve high attack performance on only 5 % poisoning ratio across safety unalignment attacks and ability degradation attacks. Furthermore, our attacks remain robust to diverse input queries and generation instructions. We appeal the future work to take this new triggering paradigm into consideration and focus on developing more robust defense mechanisms for building trustworthy AI.
Acknowledgments
This work is supported by the National Key R&D Program of China under Grant 2022YFB3103500, the National Natural Science Foundation of China under Grant 62020106013 and 62402087, the Chengdu Science and Technology Program under Grant 2023-XT00-00002-GX, the Fundamental Research Funds for Chinese Central Universities under Grant Y030232063003002, the China Postdoctoral Science Foundation under Grant BX20230060 and 2024M760356.